OpenStudy (anonymous):

Simplify: 7 SQRT 3 + 8 SQRT 3 – 2 SQRT 2 A) 13 SQRT 5 B) 13 SQRT 6 C) 15 SQRT 3 – 2 SQRT 2 I was confused if it would be B or C

OpenStudy (anonymous):

C

OpenStudy (anonymous):

C

OpenStudy (anonymous):

okay. so it looks like:\[7\sqrt{3} + 8\sqrt{3} - 2\sqrt{2}\] The first thing to do is combine like terms. Both \[7\sqrt{3}\] and \[8\sqrt{?}\] have the same radicand in common, so you can combine them to be \[15\sqrt{3}\] You have the \[-2\sqrt{?}\] left. You can't do anything to that. So, like the two above me have said, your answer will be C: \[15\sqrt{3} - 2\sqrt{2}\]

OpenStudy (anonymous):

* i meant both 7(sqrt) 3 and 8(sqrt) 3